# A Catalog of $\sim 146,000$ Radio Galaxies up to $z\sim 5$

### v1

@@ 4.2 UNIONS-VLASS Optical Source Counts
We investigate the differential counts in $r-$band for our 146,212 “clean” UNIONS-VLASS sources, as shown in Figure 10 by black circles. We also plot the optical counts of UNIONS by randomly selecting 2000 tiles in the NGP. The fall-offs at the faint end (${r}_{\mathrm{ABmag}}\geq 24$) of UNIONS-VLASS and UNIONS are highly consistent, suggesting that the limitation in associating radio galaxies with faint host galaxies is the depth of the optical survey. At the bright end (${r}_{\mathrm{ABmag}}\lesssim 17$), the optical number counts of UNIONS-VLASS sources decrease with the brighter $r-$band magnitude, whereas such a fall-off is not observed in the UNIONS. Additionally, despite the rarity of RL quasars, consistent trends are found between HSC data and HSC-FIRST RGs (Yamashita et al., 2018). Therefore, this fall-off in UNIONS-VLASS is unlikely due to saturation in the $r-$band and the lack of quasars, but arises from an elimination of the brightest nearby radio sources in “clean” VLASS2. For UNIONS-VLASS sources with $19\leq{r}_{\mathrm{ABmag}}\leq 24$, the best fit is indicated by the golden solid line. The source count has a weak dependency on the $r-$band magnitude throughout this range, despite an increasing completeness of UNIONS along with the decreasing $r-$band brightness. This implies a decreasing detectability of radio galaxies in the UNIONS catalog with an increasing apparent optical magnitude: from $0.26\%$ at $r=19$ to $0.01\%$ at $r=24$. Such a trend is also confirmed by VLASS RGs in Euclid Deep Fields-North and Fornax (reaching ${f}_{\mathrm{matches}}\sim 90\%$) down to a VIS magnitude of 28 (Zhong et al. in prep.). This could be the joint consequence of: (1) the Malmquist effect (Miley & De Breuck, 2008) subject to the detection limit of VLASS as the redshift increases; (2) the intrinsic rarity of radio AGNs at high-$z$, partially driven by the increasing energy density of the cosmic microwave background that enhances inverse Compton scattering of synchrotron-radiating electrons, which further exacerbates the selection bias (Condon & Ransom, 2016; Saxena et al., 2017).
@@
Figure 11: Left: observed radio loudness as a function of the redshift. Right: observed radio loudness as a function of the $r-$band magnitudes. These distributions include 63,019 UNIONS-VLASS radio galaxies with valid $r-$band photometric measurements. The observed radio loudness $\mathcal{R}_{\mathrm{obs}}$ is defined as the ratio of the flux density at 1.4 GHz to that of the $r-$band. The RLAGNs are characterized by $\mathcal{R}_{\mathrm{obs}}>1$ and are indicated by orange crosses, while the dark violet pluses indicate RQ AGNs. The contours in the left panel are only for RL populations. The black dots represent 3,579 radio galaxies from the HSC-FIRST catalog (Yamashita et al., 2018). We also show the detection limits of FIRST and VLASS by orange and blue solid lines, respectively.
@@
Figure 12: Left: 1.4 GHz spectral luminosity as a function of the observed radio loudness. Right: 1.4 GHz spectral luminosity as a function of the $r-$band magnitudes. These distributions include 63,019 UNIONS-VLASS radio galaxies with valid $r-$band photometric measurements. The horizontal black dashed line indicates the RL/RQAGN demarcation of ${L}_{\mathrm{1.4GHz}}\sim 10^{24}\mathrm{\ W\ Hz^{-1}}$ (Ivezić et al., 2002). The colored symbols are the same as Figure 11.
@@
Figure 13: Left: The 1.4 GHz spectral luminosity as a function of photo-z for ${\color[rgb]{0,0,0}63,019}$ UNIONS-VLASS radio galaxies (orange crosses), in comparison with HSC-FIRST AGNs (black dots) and local RLAGNs (cyan triangles). VLASS flux densities at 3 GHz are converted to those at 1.4 GHz, adopting ${\alpha}_{\mathrm{VLASS}}^{\mathrm{FIRST}}$ for those detected in FIRST and adopting $\alpha=-0.7$ for those without FIRST detection. Right: The 3 GHz spectral luminosity as a function of photo-z for UNIONS-VLASS sample, in comparison with VLA 3 GHz survey. We separate ${\color[rgb]{0,0,0}63,019}$ UNIONS-VLASS radio galaxies with valid ${z}_{\mathrm{photo}}$ into two subsamples. ${L}_{\mathrm{3GHz}}$ of those detected in both VLASS and FIRST (${\color[rgb]{0,0,0}41,803}$ in total) are calculated based on $\alpha\mathrm{{}_{VLASS}^{FIRST}}$, and of those without FIRST detection (${\color[rgb]{0,0,0}21,216}$ in total) are calculated using $\alpha=-0.7$. These sources are further divided into RL and RQ populations following the division line ${L}_{\mathrm{1.4GHz}}\sim 10^{24}\mathrm{\ W\ Hz^{-1}}$ (Ivezić et al., 2002), which is indicated by the blue solid line. In both panels, the violet line indicates the $5\sigma\ (=1.0)$ mJy detection limit of FIRST and the black one indicates the $3\sigma\ (=0.36)$ mJy detection limit of VLASS.
@@ 4.3.2 UNIONS-VLASS Subsample with Valid Photo-z
To better understand RLAGNs across cosmic time, we focus on the subsample of ${\color[rgb]{0,0,0}63,019}$ radio galaxies with valid UNIONS photo-$z$. In this subsample, 58,661 are RL and 4,358 are RQ populations based on the $\mathcal{R}_{\mathrm{obs}}$ criterion. We show the distribution of $\mathcal{R}_{\mathrm{obs}}$ as a function of the photometric redshift in the left panel of Figure 11. Compared to the HSC-FIRST catalog that has $\mathcal{R}_{\mathrm{obs}}\sim 4$ at ${z}_{\mathrm{photo}}\sim 2-4$ (Yamashita et al., 2018), the bulk of our UNIONS-VLASS radio galaxies at this redshift range have $\mathcal{R}_{\mathrm{obs}}\sim 2-3$, and there is no evolution of $\mathcal{R}_{\mathrm{obs}}$ within this redshift bin. This may be explained by the fact that the HSC-FIRST catalog uses ${i}_{\mathrm{ABmag}}$ rather than ${r}_{\mathrm{ABmag}}$, and because of the greater depth of the HSC $i-$band, Yamashita et al. (2018) captured more optically faint sources. In the panel showing $\mathcal{R}_{\mathrm{obs}}-{r}_{\mathrm{ABmag}}$ (right panel of Figure 11), we also plot the detection limit of VLASS and FIRST. We see that at ${i}_{\mathrm{ABmag}}\gtrsim 19.5$, all radio galaxies are radio-loud, and the RQ populations only exist at ${r}_{\mathrm{ABmag}}$ brighter than this magnitude, which clearly reflects the selection bias of RLAGNs based on $\mathcal{R}_{\mathrm{obs}}$, that is, this criterion favors optically faint galaxies.
@@
To investigate whether the current abundance of RLAGNs in UNIONS-VLASS is due to the faintness of the optical counterparts, we adopt an alternative approach to selecting RLAGNs via the spectral luminosity threshold ${L}_{\mathrm{radio}}\sim 10^{24}\mathrm{\ W\ Hz^{-1}}$ established from FIRST RL/RQAGNs (Ivezić et al., 2002). This threshold depends on the reduced Hubble constant and we adopted ${h}_{\mathrm{50}}$ for consistency between works, and ${L}_{\mathrm{radio}}\sim 2\times 10^{24}\mathrm{\ W\ Hz^{-1}}$ if ${h}_{\mathrm{70}}$ is used instead, leading to no significant difference. For ${\color[rgb]{0,0,0}63,019}$ UNIONS-VLASS radio galaxies with valid ${z}_{\mathrm{photo}}$, ${\color[rgb]{0,0,0}49,231}$ are RLAGNs with ${L}_{\mathrm{radio}}>10^{24}\mathrm{\ W\ Hz^{-1}}$ There are ${\color[rgb]{0,0,0}49,012}$ sources commonly identified as RLAGNs via selections based on ${L}_{\mathrm{1.4GHz}}$ and $\mathcal{R}_{\mathrm{obs}}$. Fractionally, the consistency is $\sim 99.6\%$ (49,012/49,231) for ${L}_{\mathrm{1.4GHz}}$ and $\sim 83.6\%$ (49,012/58,661) for $\mathcal{R}_{\mathrm{obs}}$, respectively.
@@
We show ${L}_{\mathrm{1.4GHz}}$ as a function of $\mathcal{R}_{\mathrm{obs}}$ and ${r}_{\mathrm{ABmag}}$ in Figure 12. It is revealed that RQAGNs classified from $\mathcal{R}_{\mathrm{obs}}<1$ generally are weak in ${L}_{\mathrm{1.4GHz}}$ in comparison with RLAGNs. Nonetheless, some RQAGNs are brighter than ${L}_{\mathrm{1.4GHz}}=10^{25}\mathrm{\ W\ Hz^{-1}}$ with ${i}_{\mathrm{ABmag}}=16-18$, suggesting that they could be RL quasars. There is a clear correlation between ${L}_{\mathrm{1.4GHz}}\ (\sim 10^{23}-10^{24}\mathrm{\ W\ Hz^{-1}})$ and $\mathcal{R}_{\mathrm{obs}}\ (\sim 0-2)$. Similarly, a weak correlation also exists for ${L}_{\mathrm{1.4GHz}}\ (\sim 10^{23}-10^{24}\mathrm{\ W\ Hz^{-1}})$ and ${r}_{\mathrm{ABmag}}\ (\sim 16-18)$. These correlations are likely associated with the cosmic evolution of $\mathcal{R}_{\mathrm{obs}}$ at $z<1$ (left panels of Figure 11 and Figure 21). On the other hand, beyond the RL-RQ demarcation of ${L}_{\mathrm{radio}}\sim 10^{24}\mathrm{\ W\ Hz^{-1}}$, the spectral luminosity has no evolution along ${r}_{\mathrm{ABmag}}\sim 22-26$.
@@
We further plot 1.4 GHz and 3 GHz spectral luminosities in the left and right panels of Figure 13, respectively. For those detected in FIRST, ${L}_{\mathrm{3GHz}}$ is calculated adopting ${\alpha}_{\mathrm{VLASS}}^{\mathrm{FIRST}}$, while those without detection adopt a spectral index of $-0.7$. We overlay the RL and RQ demarcation on the distribution of ${L}_{\mathrm{1.4GHz}}$ and scale it to that of ${L}_{\mathrm{3GHz}}$ adopting ${\alpha}_{\mathrm{3GHz}}^{\mathrm{1.4GHz}}=-0.7$. Limited by the sensitivity, both VLASS and FIRST cannot reach the depth required to detect RQAGNs at $z>1$. Therefore, our UNIONS-VLASS radio galaxies $z>1$ are (almost) radio-loud, no matter whether they are identified based on $\mathcal{R}_{\mathrm{obs}}>1$ or ${L}_{\mathrm{1.4GHz}}>10^{24}\mathrm{\ W\ Hz^{-1}}$.
@@
Figure 14: The radio color-color diagram or spectral index map of ${\alpha}_{\mathrm{FIRST=1400MHz}}^{\mathrm{LoTSS=144Mhz}}$ and ${\alpha}_{\mathrm{VLASS=3000MHz}}^{\mathrm{FIRST=1400Mhz}}$ for 30,091 UNIONS-VLASS sources with FIRST and LoTSS detections. The darker and larger marker represents a higher redshift. The spectral shapes are shown in the four quadrants, where Q1 represents upturned, Q2 represents inverted, Q3 represents power-law, Q4 represents peaked spectrum, respectively (see texts for details).
5 Discussions
@@ 5.1 Radio Color-color Diagram
The spectral shape of the synchrotron radiation and the projected linear sizes of the synchrotron-radiating regions offer implications for the evolution stage and environment of an RLAGN (Miley & De Breuck, 2008; Murgia et al., 2011; Gasperin et al., 2018; Hardcastle & Croston, 2020; Zhong et al., 2023). To characterize the spectral shapes of our UNIONS-VLASS sources, we show the radio color-color diagram (or spectral index map) between ${\alpha}_{\mathrm{FIRST=1400MHz}}^{\mathrm{LoTSS=144Mhz}}$ and ${\alpha}_{\mathrm{VLASS=3000MHz}}^{\mathrm{FIRST=1400Mhz}}$ in Figure 14, which includes 30,091 sources detected in VLASS, FIRST, and LoTSS with $\mathtt{alpha\_\{frequency1\}\_\{frequency2\}\_err}<0.2$. The expected spectral shapes covering 144 MHz to 3 GHz are indicated by the orange lines and can be divided into four quadrants (labeled Q14141-41 - 4).
Sources located in quadrant 1 have an upturned spectrum, which could be explained by a burst of synchrotron radiation subsequent to the FIRST observations conducted mostly in the 1990s, or the inhomogeneity of the radio regions observed (Condon & Ransom, 2016). Sources located in quadrant 2 have an inverted spectrum. One possible explanation is the sudden increase in the radio brightness. More commonly, it is due to the fact that synchrotron self-absorption (SSA; e.g., Tingay & de Kool, 2003; Nyland et al., 2020) dominates synchrotron radiation. Sources located in quadrant 3 have the most commonly observed power-law distribution of the non-thermal bremsstrahlung. These sources may have a steepening of the spectral index, forming the curved power-law spectrum as a result of synchrotron losses and/or inverse-Compton losses (Dermer & Menon, 2009; Blandford et al., 2019). Sources located in quadrant 4 have a peaked frequency (${\nu}_{\mathrm{p}}$) inasmuch an equilibrium between the brightness and kinetic temperatures of the synchrotron-radiating electrons (Longair, 2011). The corresponding radio galaxies could be GHz-peaked ($0.5\leq{\nu}_{\mathrm{p}}\leq 5$ GHz) or high-frequency-peaked (${\nu}_{\mathrm{p}}>5$ GHz) sources depending on their redshifts, which comprises the so-called Peaked-Spectrum (PS) sources that are young radio galaxies (see O’Dea & Saikia 2021 and references therein). Sources with inverted spectrum between 144 MHz and 3 GHz (in quadrant 2) may also have peaked/turnover frequencies above 3 GHz, making them PS sources.
@@
Figure 16: Spectral index between FIRST 1.4 GHz and VLASS 3 GHz as function of the UNIONS BPZ photo-$z$ and the projected linear size considering the deconvolution of the beam size based on VLASS observations for ${\color[rgb]{0,0,0}15,906}$ UNIONS-VLASS sources detected in FIRST and are valid in ${z}_{\mathrm{photo}}$. The colored lines are the same as Figure 15, and an additional violet line indicates the SSA-dominated inverted spectral index of $+2.5$.
@@
The sizes of the synchrotron-emitting volume matter when it comes to the classification of radio galaxies. We further show the spectral index ${\alpha}_{\mathrm{VLASS=3000MHz}}^{\mathrm{LoTSS=144MHz}}$ (${\color[rgb]{0,0,0}16,078}$ UNIONS-VLASS-LoTSS sources) and ${\alpha}_{\mathrm{VLASS=3000MHz}}^{\mathrm{FIRST=1400MHz}}$ (${\color[rgb]{0,0,0}15,906}$ UNIONS-VLASS-FIRST sources) as a function of photo-$z$ and the VLASS projected linear size of the radio regions deconvolved from the beam size in Figure 15 and 16, respectively, considering only those with reliable spectral index estimates (uncertainties smaller than 0.2). The black solid line indicates a spectral index of $-1.3$, and sources with a spectral slope steeper than this value are called Ultra-Steep-Spectrum (USS; De Breuck et al. 2000). This could be due to a significant synchrotron ageing, suggestive of well-evolved RLAGNs (van Breugel et al., 1984; Turner, 2018). USS is often used to select high-$z$ RLAGNs as energy losses could be dominated by inverse-Compton scattering, as well as being attributed to the redshifting of the frequency (Carilli & Yun, 1999; Morabito & Harwood, 2018). The blue solid line indicates a spectral index of $-0.5$. If a source has $\alpha<-0.5$ at GHz regime and its projected linear size is more compact than 20 kpc, it is then classified as Compact-Steep-Spectrum (CSS; Callingham et al. 2017; Nyland et al. 2020; Orienti et al. 2023). CSS sources are often thought to be RLAGNs at an early evolution stage, otherwise, they are actually evolved populations while their jets are confined by the ISM of their host galaxies (Murgia, 2003; Mukherjee et al., 2016). Some sources lie above ${\alpha}_{\mathrm{3GHz}}^{\mathrm{1.4GHz}}=2.5$, which is beyond the inverted spectral slope expected to be under the dominance of SSA. They are likely to be transients that undergo a sudden outburst of radio jets.
@@
In this work, we only briefly report the spectral index distribution of our UNIONS-VLASS sources. A detailed and thorough investigation of the radio classifications of our sources will be presented in the future work (Zhong et al. in preparation).
@@
Figure 17: Color-color diagrams to select dropout sources. The upper and lower panels show selections for g- ($53<z<$ 5) and r-dropout ($64<z<$ 6) sources, respectively. The black lines represent the criteria adopted to select dropouts (Eqs. (5)-(6)). The blue dots are all UNIONS-VLASS sources with valid photometry in the corresponding two-color diagram. The contours are in $1\sigma$, $2\sigma$, and $3\sigma$ confidence levels.

### v2

@@ 4.2 UNIONS-VLASS Optical Source Counts
We investigate the differential counts in $r$-band for our 146,212 “clean” UNIONS-VLASS sources, as shown in Figure 10 by black circles. We also plot the optical counts of UNIONS by randomly selecting 2000 tiles in the NGP. The fall-offs at the faint end (${r}_{\mathrm{ABmag}}\geq 24$) of UNIONS-VLASS and UNIONS are highly consistent, suggesting that the limitation in associating radio galaxies with faint host galaxies is the depth of the optical survey. At the bright end (${r}_{\mathrm{ABmag}}\lesssim 17$), the optical number counts of UNIONS-VLASS sources decrease with the brighter $r$-band magnitude. This fall-off, representing the saturation effect and intrinsic rarity of quasars (yamashita_wide_2018), is not observed in the UNIONS galaxies. Therefore, this declining trend in UNIONS-VLASS is likely to be the result of the elimination of the brightest nearby radio sources in “clean” VLASS2. For UNIONS-VLASS sources with $19\leq{r}_{\mathrm{ABmag}}\leq 24$, the best fit is indicated by the golden solid line. The source count has a weak dependency on the $r$-band magnitude throughout this range, despite an increasing completeness of UNIONS along with the decreasing $r$-band brightness. This implies a decreasing detectability of radio galaxies in the UNIONS catalog with an increasing apparent optical magnitude: from $0.26\%$ at $r=19$ to $0.01\%$ at $r=24$. This trend is also confirmed by VLASS RGs in Euclid Deep Field North and Fornax, where ${f}_{\mathrm{matches}}$ reaches $\sim 60\%$ down to a VIS magnitude of 24 (Zhong et al. in prep.). Such a decline in detectability may result from a combination of factors: (1) the Malmquist bias (miley_distant_2008), as the detection limit of VLASS increasingly excludes faint sources at higher redshifts; and (2) the intrinsic rarity of radio AGNs at high-$z$, partly due to the rising energy density of the cosmic microwave background, which enhances inverse Compton scattering of synchrotron-radiating electrons, thereby exacerbating the selection bias (condon_essential_2016; saxena_modelling_2017).
@@
Figure 11: Left: observed radio loudness as a function of the redshift. Right: observed radio loudness as a function of the $r$-band magnitudes. These distributions include 63,019 UNIONS-VLASS radio galaxies with valid $r$-band photometric measurements. The observed radio loudness $\mathcal{R}_{\mathrm{obs}}$ is defined as the ratio of the flux density at 1.4 GHz to that of the $r$-band. The RLAGNs are characterized by $\log\mathcal{R}_{\mathrm{obs}}>1$ and are indicated by orange crosses, while the dark violet pluses indicate RQ AGNs. The contours in the left panel are only for RL populations. The black dots represent 3,579 radio galaxies from the HSC-FIRST catalog (yamashita_wide_2018). We also show the detection limits of FIRST and VLASS by orange and blue solid lines, respectively.
@@
Figure 12: Left: 1.4 GHz spectral luminosity as a function of the observed radio loudness. Right: 1.4 GHz spectral luminosity as a function of the $r$-band magnitudes. These distributions include 63,019 UNIONS-VLASS radio galaxies with valid $r$-band photometric measurements. The horizontal black dashed line indicates the RL/RQAGN demarcation of ${L}_{\mathrm{1.4GHz}}\sim 10^{24}\mathrm{\ W\ Hz^{-1}}$ (ivezic_optical_2002). The colored symbols are the same as Figure 11.
@@
Figure 13: Left: The 1.4 GHz spectral luminosity as a function of photo-z for $\color[rgb]{0,0,0}63,019$} UNIONS-VLASS radio galaxies (orange crosses), in comparison with HSC-FIRST AGNs (black dots) and local RLAGNs (cyan triangles). VLASS flux densities at 3 GHz are converted to those at 1.4 GHz, adopting ${\alpha}_{\mathrm{VLASS}}^{\mathrm{FIRST}}$ for those detected in FIRST and adopting $\alpha=-0.7$ for those without FIRST detection. Right: The 3 GHz spectral luminosity as a function of photo-z for UNIONS-VLASS sample, in comparison with VLA 3 GHz survey. We separate $\color[rgb]{0,0,0}63,019$} UNIONS-VLASS radio galaxies with valid ${z}_{\mathrm{photo}}$ into two subsamples. ${L}_{\mathrm{3GHz}}$ of those detected in both VLASS and FIRST ($\color[rgb]{0,0,0}41,803$} in total) are calculated based on $\alpha\mathrm{{}_{VLASS}^{FIRST}}$, and of those without FIRST detection ($\color[rgb]{0,0,0}21,216$} in total) are calculated using $\alpha=-0.7$. These sources are further divided into RL and RQ populations following the division line ${L}_{\mathrm{1.4GHz}}\sim 10^{24}\mathrm{\ W\ Hz^{-1}}$ (ivezic_optical_2002), which is indicated by the blue solid line. In both panels, the violet line indicates the $5\sigma\ (=1.0$ mJy) detection limit of FIRST and the black one indicates the $3\sigma\ (=0.36$ mJy) detection limit of VLASS.
@@ 4.3.2 UNIONS-VLASS Subsample with Valid Photo-z
To better understand RLAGNs across cosmic time, we focus on the subsample of $\color[rgb]{0,0,0}63,019$} radio galaxies with valid UNIONS photo-$z$. In this subsample, 58,661 are RL and 4,358 are RQ populations based on the $\mathcal{R}_{\mathrm{obs}}$ criterion. We show the distribution of $\mathcal{R}_{\mathrm{obs}}$ as a function of the photometric redshift in the left panel of Figure 11. Compared to the HSC-FIRST catalog that has $\log\mathcal{R}_{\mathrm{obs}}\sim 4$ at ${z}_{\mathrm{photo}}\sim 2-4$ (yamashita_wide_2018), the bulk of our UNIONS-VLASS radio galaxies at this redshift range have $\log\mathcal{R}_{\mathrm{obs}}\sim 2-3$, and there is no evolution of $\mathcal{R}_{\mathrm{obs}}$ within this redshift bin. This may be explained by the fact that the HSC-FIRST catalog uses ${i}_{\mathrm{ABmag}}$ rather than ${r}_{\mathrm{ABmag}}$, and because of the greater depth of the HSC $i$-band, yamashita_wide_2018 captured more optically faint sources. In the panel showing $\mathcal{R}_{\mathrm{obs}}-{r}_{\mathrm{ABmag}}$ (right panel of Figure 11), we also plot the detection limit of VLASS and FIRST. We see that at ${i}_{\mathrm{ABmag}}\gtrsim 19.5$, all radio galaxies are radio-loud, and the RQ populations only exist at ${i}_{\mathrm{ABmag}}$ brighter than around 19, which clearly reflects the selection bias of RLAGNs based on $\mathcal{R}_{\mathrm{obs}}$, that is, this criterion favors optically faint galaxies.
@@
To investigate whether the current abundance of RLAGNs in UNIONS-VLASS is due to the faintness of the optical counterparts, we adopt an alternative approach to selecting RLAGNs via the spectral luminosity threshold ${L}_{\mathrm{radio}}\sim 10^{24}\mathrm{\ W\ Hz^{-1}}$ established from FIRST RL/RQAGNs (ivezic_optical_2002). This threshold depends on the reduced Hubble constant and we adopted ${h}_{\mathrm{50}}$ for consistency between works, and ${L}_{\mathrm{radio}}\sim 2\times 10^{24}\mathrm{\ W\ Hz^{-1}}$ if ${h}_{\mathrm{70}}$ is used instead, which does not result in significant differences. For $\color[rgb]{0,0,0}63,019$} UNIONS-VLASS radio galaxies with valid ${z}_{\mathrm{photo}}$, $\color[rgb]{0,0,0}49,231$} are RLAGNs with ${L}_{\mathrm{radio}}>10^{24}\mathrm{\ W\ Hz^{-1}}$ There are $\color[rgb]{0,0,0}49,012$} sources commonly identified as RLAGNs via selections based on ${L}_{\mathrm{1.4GHz}}$ and $\mathcal{R}_{\mathrm{obs}}$. Fractionally, the consistency is $\sim 99.6\%$ (49,012/49,231) for ${L}_{\mathrm{1.4GHz}}$ and $\sim 83.6\%$ (49,012/58,661) for $\mathcal{R}_{\mathrm{obs}}$, respectively.
@@
We show ${L}_{\mathrm{1.4GHz}}$ as a function of $\mathcal{R}_{\mathrm{obs}}$ and ${r}_{\mathrm{ABmag}}$ in Figure 12. It is found that RQAGNs, classified by $\log\mathcal{R}_{\mathrm{obs}}<1$, generally exhibit weaker ${L}_{\mathrm{1.4,\mathrm{GHz}}}$ values compared to RLAGNs. Nonetheless, some RQAGNs are brighter than ${L}_{\mathrm{1.4GHz}}=10^{25}\mathrm{\ W\ Hz^{-1}}$ with ${i}_{\mathrm{ABmag}}=16-18$, suggesting that they could be RL quasars. There is a clear correlation between ${L}_{\mathrm{1.4GHz}}\ (\sim 10^{23}-10^{24}\mathrm{\ W\ Hz^{-1}})$ and $\log\mathcal{R}_{\mathrm{obs}}\ (\sim 0-2)$. Similarly, a weak correlation also exists for ${L}_{\mathrm{1.4GHz}}\ (\sim 10^{23}-10^{24}\mathrm{\ W\ Hz^{-1}})$ and ${r}_{\mathrm{ABmag}}\ (\sim 16-18)$. These correlations are likely associated with the cosmic evolution of $\mathcal{R}_{\mathrm{obs}}$ at $z<1$ (left panels of Figure 11 and Figure 21). On the other hand, beyond the RL-RQ demarcation of ${L}_{\mathrm{radio}}\sim 10^{24}\mathrm{\ W\ Hz^{-1}}$, the spectral luminosity has no evolution along ${r}_{\mathrm{ABmag}}\sim 22-26$.
@@
We further plot 1.4 GHz and 3 GHz spectral luminosities in the left and right panels of Figure 13, respectively. For those detected in FIRST, ${L}_{\mathrm{3GHz}}$ is calculated adopting ${\alpha}_{\mathrm{VLASS}}^{\mathrm{FIRST}}$, while those without detection adopt a spectral index of 0.7-0.7. We overlay the RL and RQ demarcation on the distribution of ${L}_{\mathrm{1.4GHz}}$ and scale it to that of ${L}_{\mathrm{3GHz}}$ adopting ${\alpha}_{\mathrm{3GHz}}^{\mathrm{1.4GHz}}=-0.7$. Limited by the sensitivity, both VLASS and FIRST cannot reach the depth required to detect RQAGNs at $z>1$. Therefore, our UNIONS-VLASS radio galaxies $z>1$ are (almost) radio-loud, no matter whether they are identified based on $\log\mathcal{R}_{\mathrm{obs}}>1$ or ${L}_{\mathrm{1.4GHz}}>10^{24}\mathrm{\ W\ Hz^{-1}}$.
@@
Figure 14: The radio color-color diagram of ${\alpha}_{\mathrm{FIRST=1400MHz}}^{\mathrm{LoTSS=144Mhz}}$ and ${\alpha}_{\mathrm{VLASS=3000MHz}}^{\mathrm{FIRST=1400Mhz}}$ for 30,091 UNIONS-VLASS sources with FIRST and LoTSS detections. The darker and larger marker represents a higher redshift. The spectral shapes are shown in the four quadrants, where Q1 represents upturned, Q2 represents inverted, Q3 represents power-law, Q4 represents peaked spectrum, respectively (see texts for details).
5 Discussion
@@ 5.1 Radio Color-Color Diagram
[…]
Sources located in quadrant 1 have an upturned spectrum, which could be explained by a burst of synchrotron radiation subsequent to the FIRST observations conducted mostly in the 1990s, or the inhomogeneity of the radio regions observed (condon_essential_2016). Sources located in quadrant 2 have an inverted spectrum. One possible explanation is the sudden increase in the radio brightness. More commonly, it is due to the fact that synchrotron self-absorption (SSA; e.g., tingay_investigation_2003; nyland_quasars_2020) dominates synchrotron radiation. Sources located in quadrant 3 have the most commonly observed power-law distribution of the nonthermal bremsstrahlung. These sources may have a steepening of the spectral index, forming the curved power-law spectrum as a result of synchrotron losses and/or inverse-Compton losses (dermer_high_2009; blandford_relativistic_2019). Sources located in quadrant 4 have a peaked frequency (${\nu}_{\mathrm{p}}$) inasmuch an equilibrium between the brightness and kinetic temperatures of the synchrotron-radiating electrons (longair_high_2011). The corresponding radio galaxies could be GHz-peaked ($0.5\leq{\nu}_{\mathrm{p}}\leq 5$ GHz) or high-frequency peaked (${\nu}_{\mathrm{p}}>5$ GHz) sources depending on their redshifts, which comprises the so-called Peaked-Spectrum (PS) sources that are young radio galaxies (see odea_compact_2021 and references therein). Sources with inverted spectrum between 144 MHz and 3 GHz (in quadrant 2) may also have peaked/turnover frequencies above 3 GHz, making them PS sources.
@@
Figure 16: Spectral index between FIRST 1.4 GHz and VLASS 3 GHz as function of the UNIONS BPZ photo-$z$ and the projected linear size considering the deconvolution of the beam size based on VLASS observations for $\color[rgb]{0,0,0}15,906$} UNIONS-VLASS sources detected in FIRST and are valid in ${z}_{\mathrm{photo}}$. The colored lines are the same as Figure 15, and an additional violet line indicates the SSA-dominated inverted spectral index of +2.5+2.5.
@@
The sizes of the synchrotron-emitting volume matter when it comes to the classification of radio galaxies. We further show the spectral index ${\alpha}_{\mathrm{VLASS=3000MHz}}^{\mathrm{LoTSS=144MHz}}$ ($\color[rgb]{0,0,0}16,078$} UNIONS-VLASS-LoTSS sources) and ${\alpha}_{\mathrm{VLASS=3000MHz}}^{\mathrm{FIRST=1400MHz}}$ ($\color[rgb]{0,0,0}15,906$} UNIONS-VLASS-FIRST sources) as a function of photo-$z$ and the VLASS projected linear size of the radio regions deconvolved from the beam size in Figure 15 and 16, respectively, considering only those with reliable spectral index estimates (uncertainties smaller than 0.2). The black solid line indicates a spectral index of 1.3-1.3, and sources with a spectral slope steeper than this value are called ultra-steep spectrum (USS; de_breuck_sample_2000). This could be due to a significant synchrotron aging, suggestive of well-evolved RLAGNs (van_breugel_optical_1984; turner_duty-cycle_2018). USS is often used to select high-$z$ RLAGNs as energy losses could be dominated by inverse-Compton scattering, as well as being attributed to the redshifting of the frequency (carilli_radio--submillimeter_1999; morabito_investigating_2018). The blue solid line indicates a spectral index of 0.5-0.5. If a source has $\alpha<-0.5$ at GHz regime and its projected linear size is more compact than 20 kpc, it is then classified as compact steep spectrum (CSS; callingham_extragalactic_2017; nyland_quasars_2020; orienti_young_2023). CSS sources are often thought to be RLAGNs at an early evolution stage, otherwise, they are actually evolved populations while their jets are confined by the ISM of their host galaxies (murgia_spectral_2003; mukherjee_relativistic_2016). Some sources lie above ${\alpha}_{\mathrm{3GHz}}^{\mathrm{1.4GHz}}=2.5$, which is beyond the inverted spectral slope expected to be under the dominance of SSA. They are likely to be transients that undergo a sudden outburst of radio jets.
@@
In this work, we only briefly report the spectral index distribution of our UNIONS-VLASS sources. A detailed and thorough investigation of the radio classifications of our sources will be presented in the future work (Zhong et al. 2025 in preparation).
@@
Figure 17: Color-color diagrams to select dropout sources. The upper and lower panels show selections for g- ($3<z<5$) and r-dropout ($4<z<6$) sources, respectively. The black lines represent the criteria adopted to select dropouts (Eqs. (5)-(6)). The blue dots are all UNIONS-VLASS sources with valid photometry in the corresponding two-color diagram. The contours are in $1\sigma$, $2\sigma$, and $3\sigma$ confidence levels.